archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    00 and 06 00am the security isn t active on STM59 RSA59 TOU59 RON59 WAS59 SAN59 The traffic is not impacted Date Friday 31 January 2014 10 51AM Reason for closing Done Comments 3 Comment by OVH Thursday 30 January 2014 09 23AM The maintenance is over Securisation is operational again Comment by OVH Thursday 30 January 2014 16 51PM New maintenance is planned by the fibre operator on the

    Original URL path: http://status.ovh.es/?do=details&id=6211 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    Project ADSL SDSL Task Type Incident Category Backend Core Status CLOSED Percent Complete Details DSLAM down investigations underway Date Wednesday 29 January 2014 10 30AM Reason for closing Done Comments 1 Comment by OVH Wednesday 29 January 2014 10 30AM

    Original URL path: http://status.ovh.es/?do=details&id=6200 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    web service listing availability for appointment booking is down xDSL orders with line creation are affected Date Saturday 25 January 2014 01 37AM Reason for closing Done Comments 2 Comment by OVH Friday 24 January 2014 16 23PM Orange estimates

    Original URL path: http://status.ovh.es/?do=details&id=6175 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    electrical issue on the VEL78 site The devices rebooted we are investigating The service is currently available Date Friday 24 January 2014 15 10PM Reason for closing Done Comments 2 Comment by OVH Friday 24 January 2014 12 20PM We have justed opened a ticket with Orange in order to get more information on the electrical failure Comment by OVH Friday 24 January 2014 15 10PM We have received a

    Original URL path: http://status.ovh.es/?do=details&id=6173 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Modernization Category Backend Core Status CLOSED Percent Complete Details A configuration update of the TG788v2 modems was deployed late morning today It improves QoS management ont the modem and it resolves issues such as file transfers via SCP Date Wednesday

    Original URL path: http://status.ovh.es/?do=details&id=6150 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    January 2014 18 46PM Reason for closing Done Comments 4 Comment by OVH Wednesday 15 January 2014 18 34PM The modification has been delayed until January 20th Customers on the Lyon LNS de Lyon in FT collect will be moved on January 20th 00 00 to the Roubaix LNS They will be used backed to the Lyon LNS after the improvement Comment by OVH Monday 20 January 2014 17 51PM

    Original URL path: http://status.ovh.es/?do=details&id=6082 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    have been mounted on the backup LNS Comment by OVH Monday 20 January 2014 14 41PM We re replacing a card on the LNS The intervention is in progress Comment by OVH Monday 20 January 2014 16 00PM We will insert the new card into the chassis Comment by OVH Monday 20 January 2014 16 01PM unlocking sfr collect port Comment by OVH Monday 20 January 2014 16 33PM We

    Original URL path: http://status.ovh.es/?do=details&id=6071 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    the national Comment by OVH Friday 03 January 2014 01 45AM all is ok now we have changed them with no loss of bgp sessions only some ping loss Comment by OVH Tuesday 14 January 2014 21 57PM the failure has reoccurred on the routing card in collect FT national The impacted accesses are raising We are following up the failure analysis Comment by OVH Tuesday 14 January 2014 21

    Original URL path: http://status.ovh.es/?do=details&id=6047 (2015-10-15)
    Open archived version from archive



  •